OpenStudy (anonymous):

solve this equaion 7.8y+9=6.8y-2

OpenStudy (anonymous):

Isolate y terms to the left side, can you do it?

OpenStudy (anonymous):

@teresaa can you?

OpenStudy (anonymous):

so would i have to add the y

OpenStudy (anonymous):

like 7.8y+6.8y

OpenStudy (anonymous):

No, subtract 6.8y to both sides!

OpenStudy (anonymous):

so would it be 1y=9=-2

OpenStudy (anonymous):

i mean 1y+9=-2

OpenStudy (anonymous):

In the left side: 7.8 y - 6.8y = ....?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

7.8 y - 6.8y = y

OpenStudy (anonymous):

y+9= -2

OpenStudy (anonymous):

In the right side: -2 -9 = ...?

OpenStudy (anonymous):

i thought u have to divide the 2 into the 9

OpenStudy (anonymous):

so it would be -11

OpenStudy (anonymous):

so the actual answer is y= -11

OpenStudy (anonymous):

thank u'